# SALAD: Self-Adaptive Link Adaptation

### v1

@@ IV-B Over-the-air experiments
Figure 9: Knowledge distillation for learning rate $\varepsilon$ self-adaptation, with $N_{\varepsilon}=200$, $\beta=0$. The number of knots $K$ is selected via cross-validation. The SINR [ $\text{\,}\mathrm{dB}\text{/}$] evolves according to a sinusoid with time-varying frequency. The learning rate $\varepsilon$ is optimized over a grid of step 0.1. As the SINR frequency increases, $\varepsilon$ also increases, allowing SALAD to track the ground truth. Results are reproducible via [9].
@@
IV-C Knowledge distillation for learning rate self-adaptation
To conclude, we evaluate learning rate $\varepsilon$ online optimization via knowledge distillation (see Section III-A) under controlled SINR evolution in pure simulation. Every $N_{\varepsilon}$ slots, the history of ACK/NACK feedback, TBS, and selected MCS is collected. The “teacher” inference model—the piece-wise linear model (14)—is optimized using batch gradient descent, using the gradient expressions (13),(20). The number of knots $K$ is selected via cross-validation, by splitting the history into training and test sets, and minimizing the loss on the latter. Next, the “student” gradient-based inference algorithm is replayed on the same data using different candidate values of $\varepsilon$. The learning rate that minimizes the squared difference between the student’s and the teacher’s inferences is selected for use in the next slots. Figure 9 illustrates the qualitative effect of knowledge distillation when the SINR oscillates according to a sinusoid with time-varying frequency. The learning rate $\varepsilon$ adapts to the frequency of the unknown SINR, allowing SALAD to closely track the ground truth. As the frequency increases, $\varepsilon$ increases accordingly to follow faster SINR variations. Conversely, as the frequency decreases, $\varepsilon$ is reduced to prevent excessive fluctuations.
The code used to produce Figures 1, 2, 9 is available at [9].
@@ V Conclusions
In this paper, we have introduced SALAD, a novel algorithm for adaptive MCS selection for wireless communications. SALAD infers the user’s effective SINR from ACK/NACK feedback alone by minimizing a regularized cross-entropy loss with respect to predicted BLER values using gradient descent. Its learning rate can be optimized at run-time via knowledge distillation, by occasionally comparing its performance against a more powerful “teacher” inference performed in hindsight. To accelerate MCS adaptation under improving channel conditions, SALAD leverages hypothesis testing to detect SINR underestimation and safely probes high MCS values. This improves both short-term spectral efficiency and long-term SINR inference accuracy. A feedback control loop ensures that the long-term BLER remains close to the desired target.

### v2

@@ V Conclusions
In this paper, we have introduced SALAD, a novel algorithm for adaptive MCS selection for wireless communications. SALAD infers the user’s effective SINR from ACK/NACK feedback alone by minimizing a regularized cross-entropy loss with respect to predicted BLER values using gradient descent. To accelerate MCS adaptation under improving channel conditions, SALAD leverages hypothesis testing to detect SINR underestimation and safely probes high MCS values. This improves both short-term spectral efficiency and long-term SINR inference accuracy. A feedback control loop ensures that the long-term BLER remains close to the desired target.